OpenStudy (anonymous):

Help (dont really understand this) - If ln2=.693, what is the value of ln32 ? so i know theres a 5 involve in this haha becasue 2^5=32 EXCEPT i dont know how this fits to the question ...

OpenStudy (anonymous):

ohhh wait do i multiply .693 and 5 = 3.465

OpenStudy (anonymous):

the closes answer to what i got is 3.466 .... im guessing that would be the answer

OpenStudy (anonymous):

.693x32?

OpenStudy (anonymous):

mmmhhhh

OpenStudy (amistre64):

ln2=.693 ln(32) = ln(2*2*2*2*2) ln(32) = ln(2) + ln(2) + ln(2) + ln(2) + ln(2)

OpenStudy (amistre64):

or simply: ln(32) = 5 ln(2) = ln(2^5)